Find the equation of the circle whose center is at (3, -5) and whose radius is 4
1
Expert's answer
2020-05-04T19:08:24-0400

We have center of the circle "(h, k) = (3, -5)" and


Radius ="r = 4"


We know the equation of the circle with center and radius



"(x -h)^2 + (y -k)^2 = r^2"

Now we can plug the center and radius

"(x - 3)^2 + (y +5)^2 = 4^2"

"x^2 +9 -6x +y^2 +25 +10y = 16"

"x^2 + y^2 - 6x + 10y + 18 =0"

Answer: Equation of circle is "x^2 + y^2 - 6x + 10y + 18 =0"